Status Report

ISS On-Orbit Status 27 Jun 2002

By SpaceRef Editor
June 27, 2002
Filed under , ,

All ISS systems continue to function nominally,
except as noted previously or below.

After wakeup at 2:00am EDT and before breakfast, Valery Korzun and
Sergei Treschev undertook their first regular "Hematocrit"
experiment (PZE-MO-10), a biomedical blood analysis to determine the
amount of red blood cell mass.

This was followed by a session of the biomed "Sprut"
(octopus) experiment (MBI-1 5) for the cosmonauts, which serves the
study of the condition of liquids in the human body.

FE-1 Peggy Whitson’s schedule today called for her first medical EVA
fitness test, which assesses hand and finger strength with the
programmed use of a hand grip force dynamometer and a finger pinch
gauge. [The test checks the subject’s ability to perform work
while wearing spacesuit gloves rigidized by the internal suit

Later, Peggy took her first general MedOps PFE (periodic fitness
evaluation), which checks up on blood pressure and electrocardiogram
during programmed exercise on the CEVIS in the Lab.  CDR Korzun,
a qualified CMO (crew medical officer), assisted with the
examination.  Readings were taken with the BP/ECG (blood
pressure/electrocardiograph).  [BP/ECG provides automated
noninvasive systolic and diastolic blood pressure measurements while
also monitoring and displaying accurate heart rates on a continual
basis at rest and during exercise.  Whitson plugged it in the
defibrillator outlet in the Lab, since it requires USOS voltage for
operation, whereas the CHeCS HRM (Crew Health Maintenance System
heart rate monitor), which measures and records heart rate only, can
operate on 24V for use on the exercise equipment in the Russian
segment, such as the TVIS treadmill.]

FE-2 Treschev set up and then took the periodic O-OHA (on-orbit
hearing assessment) test, an EHS (environmental health systems)
examination to assess the efficacy of acoustic countermeasures.
 The two other crewmembers followed suit.  [The O-OHA
test involves minimum audibility measurements for each ear over a
wide range of frequencies and sound pressure levels, with the
crewmembers using individual-specific Prophonics earphones, Bose ANC
headsets and the SLM (sound level meter).  To conduct the
testing, the experimenter is supported by special “EarQ” software on
the MEC (medical equipment computer). The baseline test is required
for  about Flight Day 14 for the Expedition and is then
performed once per month.]

CDR Korzun and FE-2 Treschev had a two-hour period scheduled for
TORU training, in preparation for Progress M-46 (8P) docking on
Saturday morning.  They were supported by a TORU specialist at
MCC-Moscow.  [TORU (teleoperator control system) is a manual
mode through which a crewmember located in the Service Module
performs necessary guidance functions in the event of a failure of
the automated Kurs rendezvous/docking system of a Progress cargo
vehicle (TGK). In this case, the ship’s motion will be manually
controlled from the TORU control panel, using two hand controllers
and a situational display of television signals (ISS image plus
range, velocity, and relative angular position data) from a
Progress-mounted Klest-M video camera, shown on two monitors (VKU and

Sergei Treschev worked on the ZCG (Zeolite crystal growth)
experiment, first setting up the video equipment for historical
documentation, then activating the furnace by powering up EXPRESS
Rack 2 (ER2) and the IZECS (improved Zeolite electronic control
system) and enabling ER2 comm link for data.  At this point,
autoclave mixing and furnace heat-up started.  Once ZCG
operation was underway, the crew is providing  periodic systems
health monitoring. [Yesterday’s troubleshooting of the comm access
failure of the backup CPU was unsuccessful, and the ZCG is currently
operating only on the Primary CPU.]

After familiarizing herself with the procedures for nutrient,
condensate and gas sampling of the ADVASC (advanced astroculture)
experiment, Peggy Whitson collected the first samples of air from the
plant growth chamber and water from the humidity control system for
analysis on the ground.

She was also asked to obtain and record the final CSA-CP (compound
specific analyzer-combustion products) readings for all three units
today, as well as the CO2 reading of the CDMK (carbon dioxide
monitoring kit).  [If the ground finds the values nominal,
the Expedition 4 unit could be powered off for return to the ground,
and the regular O2/CO2 data requests (every Monday and Thursday) will
then only require the Exp. 5 CSA-CP for O2 and the CDMK for CO2.]

Daily routine maintenance was performed by Korzun (SOSh life
support systems servicing, IMS delta file preparation) and Whitson
(autonomous payloads status check)ZACYbr>

At 10:15 am EDT, the crew downlinked a TV message for Labor Secretary
Elaine Chao, to be presented at the Workforce Innovation 2002
Conference in Nashville, TN, on opening day, 7/9, sponsored by the
U.S. Department of Labor, Employment & Training Administration
and the National Association of Workforce Boards.  The message
was followed by a second PAO event,- a live interactive TV Q&A
exchange with WHO Radio in Des Moines, IA, via S-band (audio) and
Ku-band (video).  Both downlinks were outstanding.

Yesterday’s live PAO event with the NASA Iowa Connection was
extremely well received.  The crew answered questions from
schoolchildren about daily activities in space and gave the
participants a thrilling tour of the U.S. Lab.

The conjunction with debris of an Intercosmos rocket (object 11285)
yesterday morning (6:39am EDT) passed with no evasive action required
(which also would have impacted the rendezvous targeting of Progress

The ISS is currently flying in XPOP attitude (x-axis perpendicular to
orbit plan) for tomorrow’s testing of the Kurs system with Progress
8P.  [For the radio transmission tests, the cargo ship will
be trailing the ISS slightly above its (negative) velocity vector
"bar" at a range of approximately 30 km (and increasing)
while the station, in its XPOP attitude, "rotates" with
respect to it.]

Today’s target areas for the U.S. CEO program were W.
Mediterranean Dust and Smog
(a weak Atlantic storm off the
coast of Portugal had the potential to generate dust plumes over
northwestern Africa this pass.  As ISS crossed Morocco and then
the coast of Algeria, the crew was to look to the left of track in
oblique views for evidence of dust clouds streaming northward)
E. Mediterranean Dust and Smog (on this descending pass,
crew was to look for smog moving off the Balkan Peninsula over the
Aegean Sea and then for more near the coast of southwestern Turkey,
using oblique views to the right of track. Sun glint views are
possible and useful)
, Gulf of St. Lawrence (with fair
weather over Labrador this pass, as the crew approached northern
Newfoundland they were look well to the left of track towards the
coast of Labrador for any remaining sea ice or icebergs)
Canadian Rocky Mountains (this was first pass today over
this target.  It offered excellent, near-nadir views of most of
the features of interest here. Use the long lenses of the ESC
[electronic still camera] was advised, to carefully map the details
of the small ice fields and glaciers here)
, Eastern United
(Southern New England was in the vicinity of a
stationary front.  Satellite imagery and surface reports
suggested heavy aerosol concentrations to the right of track.
 Of interest: oblique and limb views.  The second pass over
this target was best suited for more oblique context photos to the
left of track, for documenting the snow pack and complex topographic
features of this target)

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 2:13 pm

Environmental Control and Life Support (ECLSS) and
Thermal Control (TCS):
Elektron O2 generator is powered On (24-amp mode). Vozdukh
CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).
 U.S. CDRA CO2 scrubber is Off (failed).  BMP Harmful
Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify

SM Working Compartment:  Pressure (mmHg) — 758,
temperature (deg C) — 27.2, ppO2 (mmHg) — 141.6, ppCO2 (mmHg) —
SM Transfer Compartment:  Pressure (mmHg) — 762,
temperature (deg C) — 21.2.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C)
— 21.3.
Node:  Pressure (mmHg) — 752.52, temperature (deg C) —
25.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 753.6, temperature (deg C)
— 24.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 753.5,
temperature (deg C) — 24.3; shell heater temp (deg C) — 24.3, ppO2
(mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 26.2
PMA-2:  Shell heater temp (deg C) — 12.1

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not
available because MCA [major constituent analyzer] is failed and in
Extended Life mode
[= a state that preserves mass spectrometer
vacuum but produces no pp data]). MSA (mass spectrometer assembly)
and VGA (verification gas assembly) have been removed for return to

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B in
Autotrack mode.
SM batteries: Battery #1 reported failed (6/21); battery #8 in
"Cycle" mode. All other batteries (6) are in "Partial
charge" mode.
FGB: Battery #2 is offline (ROM mode); battery #5 in
"Cycle" mode.  All other batteries (4) are in “Partial
charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1
is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Attitude Source:
State vector — US GPS (SIGI string 1)
Attitude — Russian
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is powered Off (due
to consistently reaching upper operational temperature during

SSRMS/Canadarm2 at MBS PDGF (mobile base system/power &
data grapple fixture), with Keep Alive power on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of last night, 10:27pm EDT [= epoch]):
Mean altitude — 392.6 km
Apogee — 397.6 km
Perigee — 387.7 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.00073
Orbits per 24-hr. day — 15.58
Altitude decrease — 200 m (mean) in last 24 hours
Solar Beta Angle —  -28.1 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. ’98) — 20557
Current Flight Attitude — XPOP (x-axis perpendicular to orbit
plane = “sun-fixed” [yaw: -1.2 deg, pitch: -7.4 deg., roll: 0

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.